OpenStudy (anonymous):

Help last question!! Explain he steps you would use to solve the equation. Find the solution. 8^2x=4096

OpenStudy (anonymous):

I have no Idea

OpenStudy (anonymous):

Ok, thank you for trying @116601

OpenStudy (anonymous):

@mcnabkt

OpenStudy (lizz123):

Write your problem (numbers) and scroll down and it gives you detailed information

OpenStudy (lizz123):

@feathert03s

OpenStudy (anonymous):

@lizz123 i want the x to be up with the 2 how do i do that

OpenStudy (lizz123):

you write the 2 with the x it does numbers and letters

OpenStudy (anonymous):

This is what its doing i need it to have 2x as an exponent

OpenStudy (anonymous):

@lizz123

OpenStudy (lizz123):

oh idk sorry

OpenStudy (mathmate):

Hint: Express both sides as a power of 8. \(\Large 8^{2x}=4096=8^4\) Solve for x.

OpenStudy (anonymous):

what? that doesn't make sense

OpenStudy (anonymous):

where did you get ^4 from @mathmate

OpenStudy (anonymous):

@calculusfunctions

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (mathmate):

If you keep dividing 4096 by eight, and count how many times you divide before you get a one. I.e. 4096/8/8.../8 will tell you 8*8*..*8 = 8 raised to how many times will give 4096.

OpenStudy (mathmate):

The other way to do it is to factorize 4096 into powers of two, and work from there.

OpenStudy (anonymous):

ohk so x=?

OpenStudy (calculusfunctions):

Try to equate the bases (which you can in your question). In other words, first write 4096 as a power of 8.

OpenStudy (anonymous):

we have 8^4

OpenStudy (mathmate):

So that's the right hand side of 8^4. Can you now solve for x?

OpenStudy (anonymous):

no because there are two =

OpenStudy (calculusfunctions):

Since the 4th power of 8 is 4096,\[8^{2x}=4096\]is equivalent to\[8^{2x}=8^{4}\]Now if the bases are equal, then the exponents are equal. Thus\[2x =4\]

OpenStudy (calculusfunctions):

Understand @feathert03s ?

OpenStudy (mathmate):

The middle is just an intermediate step. The left hand side still equals the right hand side, right?

OpenStudy (anonymous):

ok, so i put 8^22 = 4096 = 8^4 8^2x = 8^4 2x= 4

OpenStudy (calculusfunctions):

Yes!

OpenStudy (calculusfunctions):

Now solve for x.

OpenStudy (mathmate):

Absolutely! \(\color{blue}{\Large\checkmark}{Correct!}\)

OpenStudy (anonymous):

x=2

OpenStudy (calculusfunctions):

You got it!

OpenStudy (anonymous):

sweet thank you both!!

OpenStudy (mathmate):

You're welcome! :)

OpenStudy (calculusfunctions):

Welcome!